OpenStudy (anonymous):

find the slope of the line passing through the points a(-8 -6) and b(7 0).

OpenStudy (anonymous):

Slope of the line AB: (yB - yA)/(xB - xA) = (0 + 6)/(7 + 8) =

OpenStudy (anonymous):

0.4?

OpenStudy (anonymous):

Correct